Assume that your group represents the Credit Manager of a North Vancouver Credit Union and that Mr. Wayne Gretski, on his way please do C part only
Answer #1
PAYMENT FOR FIRST 5 YEARS
Monthly interest rate =r
Semi annual interest rate=(4.3/2)%= 0.0215
(1+r)^6=1.0215
1+r=1.0215^(1/6)= 1.003551648
Rate Monthly interest rate =r= 0.003551648
Pv Amount of loan=760000-295000 $465,000
Nper Number of months of mortgage 300 (25*12)
PMT Monthly payment $2,522.20
(Using PMT function of excel)
Loan Balance at the end of 60 months =Present Value of future payments for(300-60)=240 months
Rate Monthly interest rate =r= 0.003551648
Nper Number of months of future payment 240
Pmt Monthly payment $2,522.20
PV Loan Balance at the end of 60 months $406,890
(Using PV function of excel)
Lump sum payment at end of 5 years $115,000
PAYMENT FOR NEXT 5 YEARS
Pv1 Loan Balance after lump sum payment $291,890 (406890-115000)
Semi annual interest rate =(4.75/2)% 0.02375
Rate 1 Monthly interest rate =(1.02375^(1/6))-1 0.004705508
Nper1 Number of months of Mortgage=20*12 240
PMT1 Monthly payment $2,032.12
(Using PMT function of excel)
Loan Balance at the end of 120 months =Present Value of future payments for(240-60)=180months
Rate Monthly interest rate =r= 0.004705508
Nper Number of months of future payment 180
Pmt Monthly payment $2,032.12
PV2 Loan Balance at the end of 120 months $246,352
(Using PV function of excel)
C.(i) Size of Principal balance being refinaced $246,352
(ii) NEW MONTHLY PAYMENT
Pv2 Loan Balance $246,352
Rate 2 Monthly interest rate =(5.25/12)% 0.4375%
Nper2 Number of months of Mortgage=15*12 180
PMT2 Monthly payment $1,980.37
(Using PMT function of excel)
NEW MONTHLY PAYMENT $1,980.37
ANNUAL BALANCE:
First 5 Years
Principal Payment using PPMT function with Rate =0.003551648,Nper=300,Per=Month,Pv=-465000
Per PPMT
Month Principal Payment Principal Balance End of Year Principal Balance
0 $465,000 0 $465,000
1 $870.69 $464,129.31 1 $454,345
2 $873.78 $463,255.54 2 $443,227
3 $876.88 $462,378.66 3 $431,626
4 $880.00 $461,498.66 4 $418,489
5 $883.12 $460,615.54 5 $291,890
6 $886.26 $459,729.28 6 $283,779
E17 =PV(E14,E15,-E16) A B E F G H I J K 0.0215 C 1 PAYMENT FOR FIRST 5 YEARS Monthly interest rate =r Semi annual interest raE32 x fc =PV(E29,E30,-E31) E F G H I J 12 13 14 Rate 15 Nper 16 Pmt 17 PV 18 19 Loan Balance at the end of 60 months =PresentE10 x fc =PMT(E7,99,-E8) F 0.0215 PAYMENT FOR FIRST 5 YEARS Monthly interest rate =r Semi annual interest rate=(4.3/2)%= (1+rE25 X Fax =PMT(E23, E24,-E21) B F I G H C Pmt PV E $2,522.20 $406,890 $115,000 Pv1 Monthly payment Loan Balance at the end ofE40 x fc =PMT(E38, E39,-E37) H F G (406890-115000) BC D Pv1 Loan Balance after lump sum payment Semi annual interest rate =(4File Home Insert Page Layout Formulas Data Review View H elp Tell me what you want to do F51 X for =PPMT(0.003551648, E51,300Per First 5 Years Principal Payment using PPMT function with Rate =0.003551648,Nper=300, Per=Month, Pv=-465001 PPMT PrincipalX axis: End of Year, Yaxis: Principal balance $500,000 $450,000 $400,000 $350,000 $300,000 T$250,000 $200,000 $150,000 $100,0First 5 Years Per Month PPMT Principal Principal Payment Balance $465,000 1 $870.69 $464,129.31 2 $873.78 $463,255.54 3 $876.NEXT 5 YEARS Principal Payment using PPMT function with Rate =0.004705508,Nper=240,Per=Month,Pv=-291890 Per PPMT Principal PrLAST 15 YEARS Principal Payment using PPMT function with Rate =0.4375%,Nper=180, Per=Month, Pv=-246352 Per PPMT Principal Pri